# A Chung-Fuchs type theorem for skew product dynamical systems

### v1

@@ 1. Introduction
    is an i.i.d. sequence of random variables under $\mathbb{Q}_{x}$ such that there exist two disjoint intervals $I_{x},J_{x}\subset(0,\infty)$ such that $\mathbb{Q}_{x}(F(x,\cdot)\in I_{x})\cdot\mathbb{Q}_{x}(F(x,\cdot)\in J_{x})>0$.
For $n\geq 1$ denote by
$S_{n}F=F+F\circ\hat{T}+\cdots+F\circ\hat{T}^{n-1}.$
Let $f:X\to\mathbb{R}$ be a measurable function on $X$. One may view $f$ as a measurable function on $X\times\Omega$ by setting $f(x,\omega)=f(x)$. We have the following extension of Chung and Fuchs theorem.
Theorem 1.3.
Assume (A). If $\mu$ is ergodic and
$\mu(\lim_{n\to\infty}\frac{S_{n}f}{n}=0)=\mathbb{Q}(\lim_{n\to\infty}\frac{S_{n}F}{n}=0)=1,$
then
$\mathbb{Q}(\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty)>0.$
Remark 1.4.
If we further assume that $\mathbb{Q}$ is $\hat{T}$-ergodic, then, since $\{\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty\}$ is a $\hat{T}$-invariant set, Theorem 1.3 implies $\mathbb{Q}({\displaystyle\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty})=1$. If we also assume $-F$ satisfies (A) then $\mathbb{Q}({\displaystyle\liminf_{n\to\infty}S_{n}f+S_{n}F=-\infty})=1$.
Remark 1.5.
Assumption (A) holds for a more general class of coupling measures $\mathbb{Q}$ than the Peyrière measure studied in [3]. In particular we are not requiring that the law of $F(x,\cdot)$ under $\mathbb{Q}_{x}$ to be independent of $x$.
Remark 1.6.
In the proofs we added more details comparing to [3]. First we provided a more detailed proof of Dekking’s recurrence theorem for our need. Then we proved Lemma 2.1 to ensure that under the recurrence assumption a “suspension flow” induced from an infinite dynamical system is still a m.p.d.s.. We also added an induced system argument to deduce that a measurable (not necessarily integrable) coboundary $g=f-f\circ T$ with $f,g\geq 0$ is equal to zero almost everywhere.
2. Proofs
2.1. Proof of Theorem 1.2
Proof.
For $c>0$ and $x\in X$ define
$\tau_{c}^{0}(x)=\inf\{n\geq 1:|S_{n}f(x)|\leq c\},$
and for $k\geq 1$ define
$\tau_{c}^{k}(x)=\inf\{n\geq\tau_{c}^{k-1}(x)+1:|S_{n}f(x)|\leq c\}.$
It is equivalent to prove that for all $k\geq 0$ and $c>0$, $\tau_{c}^{k}(x)<\infty$ for $\mu$-a.e. $x\in X$.
Set $\tau^{-1}_{c}(x)=0$. Hence for $k=0$, $\tau^{k-1}_{c}(x)<\infty$ for all $c>0$ and $\mu$-a.e. $x\in X$.
We shall prove the statement by induction. Assume that for all $c>0$, $\tau^{k-1}_{c}(x)<\infty$ for $\mu$-a.e. $x\in X$. Fix $a,b,c>0$ such that $a+b=c$. For $n\geq 1$ denote by
$R_{n}(x)=\lambda\Big{(}\bigcup_{l=1}^{n}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x))\Big{)},$
where $\lambda$ is the Lebesgue measure on $\mathbb{R}$ and $I_{a}(u)=[u-a/2,u+a/2]$ stands for the interval of length $a$ centered at $u\in\mathbb{R}$. By the transition invariance of Lebesgue measure we have that $R_{n}$ is a subadditive function:
$R_{n+m}(x)\leq$ $\lambda\Big{(}\bigcup_{l=1}^{n}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x))\Big{)}+\lambda\Big{(}\bigcup_{l=n+1}^{n+m}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x))\Big{)}$
$=$ $\lambda\Big{(}\bigcup_{l=1}^{n}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x))\Big{)}+\lambda\Big{(}\bigcup_{l=n+1}^{n+m}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x)-S_{\tau^{k-1}_{b}(x)+n}f(x))\Big{)}$
$=$ $R_{n}(x)+\lambda\Big{(}\bigcup_{l=1}^{m}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(T^{n}(x)))\Big{)}$
$=$ $R_{n}(x)+R_{m}(T^{n}x).$
Thus by Kingman’s subadditive ergodic theorem we have that for $\mu$-almost every $x\in X$,
$\lim_{n\to\infty}\frac{R_{n}(x)}{n}=\inf_{n\geq 1}\frac{\mathbb{E}_{\mu}(R_{n})}{n}.$
Using again the transition invariance of Lebesgue measure we get that
$R_{n+1}(x)=$ $\lambda\Big{(}\bigcup_{l=1}^{n+1}I_{a}(S_{\tau^{k-1}_{b}f(x)+l}f(x))\Big{)}$
$=$ $\lambda\Big{(}I_{a}(S_{\tau_{b}^{k-1}(x)+1}(x))\cup\bigcup_{l=1}^{n}I_{a}(S_{\tau_{b}^{k-1}(x)+1}f(x)+S_{l}f(T^{\tau_{b}^{k-1}(x)+1}(x)))\Big{)}$
$=$ $\lambda\Big{(}I_{a}(0)\cup\bigcup_{l=1}^{n}I_{a}(S_{l}f(T^{\tau^{k-1}_{b}(x)+1}(x)))\Big{)}$
$=$ $\lambda\Big{(}I_{a}(0)\cup\bigcup_{l=1}^{n}I_{a}(S_{l}f(T_{b,k}(T(x))))\Big{)},$
where we denote by
$T_{b,k}(x)=T^{\tau^{k-1}_{b}(x)}(x).$
By the $T$-invariance of $\mu$ and Fubini’s theorem we deduce that
$\mathbb{E}_{\mu}(R_{n+1})-\mathbb{E}_{\mu}(R_{n})$
$=$ $\int_{X}\lambda\Big{(}I_{a}(0)\cup\bigcup_{l=1}^{n}I_{a}(S_{l}f(T_{b,k}(x)))\Big{)}-\lambda\Big{(}\bigcup_{l=1}^{n}I_{a}(S_{l}f(T_{b,k}(x)))\Big{)}\,\mu(\mathrm{d}x)$
$=$ $\int_{X}\lambda\Big{(}I_{a}(0)\setminus\bigcup_{l=1}^{n}I_{a}(S_{l}f(T_{b,k}(x)))\Big{)}\,\mu(\mathrm{d}x)$
$=$ $\int_{X}\int_{I_{a}(0)}\prod_{l=1}^{n}\mathbf{1}_{\{y\not\in I_{a}(S_{l}f(T_{b,k}(x)))\}}\,\lambda(\mathrm{d}y)\,\mu(\mathrm{d}x)$
$=$ $\int_{X}\int_{I_{a}(0)}\prod_{l=1}^{n}\mathbf{1}_{\{S_{l}f(T_{b,k}(x))\not\in I_{a}(y)\}}\,\lambda(\mathrm{d}y)\,\mu(\mathrm{d}x)$
$=$ $\int_{I_{a}(0)}\mu\big{(}x:S_{1}f(T_{b,k}(x))\not\in I_{a}(y),\cdots,S_{n}f(T_{b,k}(x))\not\in I_{a}(y)\big{)}\,\lambda(\mathrm{d}y).$
This means that $\mathbb{E}_{\mu}(R_{n+1})-\mathbb{E}_{\mu}(R_{n})$ decreasingly converges to
$\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}(x))\not\in I_{a}(y)\text{ for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y)$
as $n\to\infty$. Therefore
$\inf_{n\geq 1}\frac{\mathbb{E}_{\mu}(R_{n})}{n}=$ $\inf_{n\geq 1}\frac{\sum_{m=0}^{n-1}\mathbb{E}_{\mu}(R_{m+1})-\mathbb{E}_{\mu}(R_{m})}{n}$
$=$ $\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}(x))\not\in I(y)\text{ for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y).$
Thus for $\mu$-almost every $x\in X$,
$\lim_{n\to\infty}\frac{R_{n}(x)}{n}=\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}(x))\not\in I_{a}(y)\text{ for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y).$
By the assumption that $\lim_{n\to\infty}\frac{1}{n}S_{n}f(x)=0$ and $\tau^{k-1}_{b}(x)<\infty$ for $\mu$-a.e. $x\in X$, we get that for $\mu$-almost every $x\in X$,
$\lim_{n\to\infty}\frac{S_{\tau^{k-1}_{b}(x)+n}f(x)}{n}=$ $\lim_{n\to\infty}\frac{S_{\tau^{k-1}_{b}(x)+n}f(x)}{\tau^{k-1}_{b}(x)+n}$
$=$ $0$
Thus for any $\epsilon>0$ there exists $n_{\epsilon}(x)$ such that for all $n\geq n_{\epsilon}(x)$, $S_{\tau^{k-1}_{b}(x)+n}f(x)\in[-\epsilon n,\epsilon n]$. This means that for $n\geq n_{\epsilon}(x)$,
$R_{n}(x)\leq R_{n_{\epsilon}(x)}(x)+\lambda([-\epsilon n-a,\epsilon n+a])=R_{n_{\epsilon}(x)}(x)+2\epsilon n+2a.$
Hence
$\lim_{n\to\infty}\frac{R_{n}(x)}{n}\leq 2\epsilon.$
Since $\epsilon$ is arbitrary, we get that for $\mu$-almost every $x\in X$,
$\lim_{n\to\infty}\frac{R_{n}(x)}{n}=0.$
This implies that for $\mu$-almost every $x\in X$,
$\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}(x))\not\in I_{a}(y)\text{ for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y)=0.$
Note that for $y\in I_{a}(0)=[-a/2,a/2]$ we have $I_{a}(y)=[y-a/2,y+a/2]\subset[-a,a]$. Furthermore, by the definition of $\tau_{b}^{k-1}(x)$ we have $S_{\tau_{b}^{k-1}(x)}f(x)\in[-b,b]$. Then by
$S_{\tau^{k-1}_{b}(x)+l}f(x)=S_{\tau^{k-1}_{b}(x)}f(x)+S_{l}f(T_{b,k}x),$
where when $k=0$ we may use the convention that $S_{0}f(x)=0$, we deduce that
$0=$ $\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}x)\not\in I_{a}(y)\text{for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y)$
$\geq$ $\lambda(I_{a}(0))\cdot\mu\big{(}x:S_{l}f(T_{b,k}x)\not\in[-a,a]\text{ for all }l=1,2,\cdots\big{)}$
$\geq$ $a\cdot\mu\big{(}x:S_{n}f(x)\not\in[-a-b,a+b]\text{ for all }n\geq\tau^{k-1}_{b}(x)+1\big{)}$
$=$ $a\cdot\mu\big{(}x:S_{n}f(x)\not\in[-c,c]\text{ for all }n\geq\tau^{k-1}_{b}(x)+1\big{)}$
$\geq$ $a\cdot\mu\big{(}x:S_{n}f(x)\not\in[-c,c]\text{ for all }n\geq\tau^{k-1}_{c}(x)+1\big{)}.$
This implies that $\mu\big{(}x:S_{n}f(x)\not\in[-c,c]\text{ for }n\geq\tau^{k-1}_{c}(x)+1\big{)}=0$, or in other words, for $\mu$-almost every $x\in X$,
$\tau_{c}^{k}(x)=\inf\{n\geq\tau^{k-1}_{c}(x)+1:S_{n}f(x)\in[-c,c]\}<\infty.$
2.2. Proof of Theorem 1.3
Consider the transformation $T_{f}$ on $X\times\mathbb{R}$:
$T_{f}(x,t)=(Tx,t+f(x)).$
Recall that $\lambda$ denotes the Lebesgue measure on $\mathbb{R}$. It is easy to verify that the product measure $\theta=\mu\times\lambda$ is $T_{f}$-invariant, thus $(X\times\mathbb{R},\mathcal{B}\otimes\mathcal{B}(\mathbb{R}),T_{f},\theta)$ forms a measure-preserving dynamical system. But the measure $\theta$ is a $\sigma$-finite measure with infinite mean, hence we need to induce this system to a finite one. In order to do so we need the system to be conservative. By Theorem 1.2 this is granted by (1.3).
Lemma 2.1.
If (1.3) holds and $\mu$ is ergodic then $T_{f}$ is conservative.
Proof.
It is enough to show that for set $B\in\mathcal{B}$ with $\mu(B)>0$ and $(a,b)\subset\mathbb{R}$, for every $n\geq 1$ there exists $m\geq n$ such that, writing $A=B\times(a,b)$,
$\theta(A\cap T_{f}^{-m}A)>0.$
Fix $B\in\mathcal{B}$ with $\mu(B)>0$. The induced system $(B,\mathcal{B}|_{B},T^{B},\mu_{B})$ is measure-preserving and ergodic (see [7, Proposition 1.4.1] for example), where $\mu_{B}=\frac{1}{\mu(B)}\mu|_{B}$ and $T^{B}(x)=T^{\rho_{B}(x)}(x)$ for $x\in B$, where
$\rho_{B}(x)=\inf\{n\geq 1:T^{n}(x)\in B\}$
is the first return time of $x\in B$. By Kac formula (see [7, Theorem 1.2.2] for example) we have that $\rho_{B}$ is integrable and $\int_{B}\rho_{B}\,\mathrm{d}\mu\leq 1$. Write $S_{n}^{B}\rho_{B}(x)=\rho_{B}(x)+\cdots+\rho_{B}\circ(T^{B})^{n-1}(x)$. Then Birkhoff ergodic theorem implies that
(2.1) $\lim_{n\to\infty}\frac{S_{n}^{B}\rho_{B}(x)}{n}=\tau(B):=\int_{X}\rho_{B}(x)\,\mu_{B}(\mathrm{d}x)$
exists for $\mu_{B}$-a.e. $x\in B$ and $\tau(B)\leq\frac{1}{\mu(B)}<\infty$.
For $x\in B$ write $f^{\rho_{B}}(x)=f(x)+\cdots+f\circ T^{\rho_{B}(x)-1}(x)$ and
$S_{n}^{B}f^{\rho_{B}}(x)=f^{\rho_{B}}(x)+\cdots+f^{\rho_{B}}\circ(T^{B})^{n-1}(x).$
Then (1.3) and (2.1) implies that for $\mu_{B}$-a.e. $x\in B$,
$\lim_{n\to\infty}\frac{S_{n}^{B}f^{\rho_{B}}(x)}{n}=\lim_{n\to\infty}\frac{S_{n}^{B}f^{\rho_{B}}(x)}{S_{n}^{B}\rho_{B}(x)}\frac{S_{n}^{B}\rho_{B}(x)}{n}=0.$
In other words, the measurable function $f^{\rho_{B}}$ satisfies (1.3) in terms of the induced system $(B,\mathcal{B}|_{B},T^{B},\mu_{B})$. Hence by Theorem 1.2 the generalised random walk $\{S_{n}^{B}f^{\rho_{B}}\}_{n\geq 1}$ is recurrent in this induced system.
Take $c>0$ such that $a+c<b-c$. For $(x,t)\in B\times(a+c,b-c)$ and $n\geq 1$, by recurrence for $\mu_{B}$-a.e. $x\in B$ one can find $k(x)\geq 1$ such that $S_{k(x)}^{B}\rho_{B}(x)\geq n$ and $S_{k(x)}^{B}f^{\rho_{B}}(x)\in(-c,c)$. Since $\mu(B)>0$, there exists $m\geq n+1$ such that $\mu_{B}(S_{k(x)}^{B}\rho_{B}(x)=m)>0$. Therefore
$\theta(A\cap T_{f}^{-m}A)$ $\,\geq\theta((x,t)\in B\times(a+c,b-c):S_{k(x)}^{B}\rho_{B}(x)=m)$
$\,\geq\mu(x\in B:S_{k(x)}^{B}\rho_{B}(x)=m)\cdot(b-a-2c)>0.$
Now fix $E=X\times[-c,c]$ for some $c>0$. By conservativeness (Lemma 2.1) the induced system $(E,\mathcal{B}_{E},T_{f}^{E},\theta_{E})$ is a m.p.d.s. with $\theta_{E}(E)=1$ (see [1, Proposition 1.5.3] for example). The measure $\theta_{E}$ takes the form
$\theta_{E}=\frac{1}{\theta(E)}\theta\times\lambda|_{E}=\mu\times\lambda_{c},$
where we denote by $\lambda_{c}=\frac{1}{2c}\lambda|_{[-c,c]}$ the normalised Lebesgue measure on $[-c,c]$. We may now define the following induced skew-product dynamical system
$(\overline{X},\overline{\mathcal{B}},\overline{T},\overline{\mathbb{Q}}),$
@@ 2. Proofs
    $\overline{X}=X\times\Omega\times[-c,c]$ and $\overline{B}=\mathcal{B}\otimes\mathcal{A}\times\mathcal{B}([-c,c])$.
    $\overline{T}(x,\omega,t)=(\hat{T}^{\rho_{E}(x,t)}(x,\omega),t+f(x)+\cdots+f\circ T^{\rho_{E}(x,t)-1}(x))$.
    ¯=×$\overline{\mathbb{Q}}=\mathbb{Q}\times\lambda_{c}$.
It is easy to verify that $\overline{\mathbb{Q}}$ is $\overline{T}$-invariant. Note that the projection of $\overline{\mathbb{Q}}$ onto $E=X\times[-c,c]$ is $\theta_{E}$ and the disintegration of $\overline{\mathbb{Q}}$ with respect to $\theta_{E}$ is ¯$\overline{\mathbb{Q}}_{x,t}=\mathbb{Q}_{x}$ for $\theta_{E}$-a.e. $(x,t)\in E$.
Proof of Theorem 1.3.
We shall use the filling scheme [6]. For $g\in\mathbb{R}$ denote by $g^{+}=\max\{g,0\}$ and $g^{-}=\max\{-g,0\}$.
We may view both $f$ and $F$ as functions on $\overline{X}$ by taking $f(x,\omega,t)=f(x)$ and $F(x,\omega,t)=F(x,\omega)$. For $n\geq 1$ define
$G_{n}:=\max_{1\leq k\leq n}\overline{S}_{k}F,$
where we denote by
$\overline{S}_{k}F=F+F\circ\overline{T}+\cdots+F\circ\overline{T}^{k-1}.$
Then we have
$G_{n+1}^{+}-G_{n}^{+}\circ\overline{T}=F+G_{n+1}^{-}.$
Let $G=\lim_{n\to\infty}G_{n}$. Assume that ¯$\overline{\mathbb{Q}}(G<\infty)=1$. Then we may write
$G^{+}-G^{+}\circ\overline{T}=F+G^{-}.$
Let
$\kappa:=\inf\{n\geq 1:F\circ\overline{T}^{n-1}>0\}.$
Then, since for all $n\geq 1$, $G_{n}\circ\overline{T}^{\kappa-1}\geq F\circ\overline{T}^{\kappa-1}>0$, we get that $G\circ\overline{T}^{\kappa-1}\geq F\circ\overline{T}^{\kappa-1}>0$ hence $G^{+}\circ\overline{T}^{\kappa-1}=G\circ\overline{T}^{\kappa-1}$ and $G^{-}\circ\overline{T}^{\kappa-1}=0$. Denote by $H(x,\omega,t)=G^{\kappa(x,\omega,t)-1}(x,\omega,t)$ and $h(x,\omega,t)=F^{\kappa(x,\omega,t)-1}(x,\omega,t)$. Then we deduce that
$H-H\circ\overline{T}=h.$
For $u\in\mathbb{R}$ denote by
$\Phi_{u}(x):=\mathbb{E}_{\overline{\mathbb{Q}}_{x,t}}(e^{iuH(x,\cdot,t)})$
and
$\phi_{u}(x):=\mathbb{E}_{\overline{\mathbb{Q}}_{x,t}}(e^{iuh(x,\cdot,t)}).$
Then by (A) we have
$\Phi_{u}(x,t)=\phi_{u}(x,t)\cdot\Phi_{u}\circ T_{f}^{E}(x,t).$
Write $f(x,t)=-\log|\Phi_{u}(x,t)|$ and $g(x,t)=-\log|\phi_{u}(x,t)|$. Then we have that $f,g\geq 0$ and
$f-f\circ T_{f}^{E}=g.$
Take $N>0$ large enough such that
$E_{N}=E\cap\{f\leq N\}$
has positive $\theta_{E}$-mass. Let $(E_{N},\mathcal{B}_{N},T_{N},\theta_{N})$ denote the induced dynamical system by $E_{N}$. We get $f-f\circ T_{N}=g_{N}$, where
$g_{N}=g+g\circ T_{f}^{E}+\cdots+g\circ(T_{f}^{E})^{\rho_{E_{N}}(x,t)-1}.$
@@
Now on $E_{N}$ we have that $0\leq f\leq N$, $0\leq f\circ T_{N}\leq N$ and $0\leq g_{N}\leq f\leq N$ are all bounded functions hence integrable, therefore by the $T_{N}$-invariance of $\mu_{N}$ and Birkhorff ergodic theorem,
$\mathbb{E}_{\theta_{N}}(g_{N}\,|\,\mathcal{I}_{N})(x,t)=0$
@@
for $\theta_{N}$-a.e. $(x,t)\in E_{N}$, where $\mathcal{I}_{N}$ is the $T_{N}$-invariant $\sigma$-algebra of $E_{N}$. Since $g$ is non-negative function, $g_{N}$ is also nonnegative, we deduce that $g_{N}(x,t)=0$ and therefore $g(x,t)=0$ for $\theta_{N}$-a.e. $(x,t)\in E_{N}$. Since $N$ is arbitrary, we deduce that $g(x,t)=0$ or equivalently $|\phi_{u}(x,t)|=1$ for $\theta$-a.e. $(x,t)\in E$. This implies that for $\theta_{E}$-a.e. $(x,t)\in E$, $h(x,\cdot,t)=F^{\kappa(x,\cdot,t)-1}(x,\cdot,t)$ is $\mathbb{Q}_{x,t}$-a.s. a constant. But this is not possible under (A), the reason is the following: Recall the disjoint intervals $I_{x}$ and $J_{x}$ from (A). We may choose the intervals a bit smaller if necessary so that $a_{x}=\inf I_{x}\cup J_{x}>0$. Then we have $\mathbb{Q}_{x}(F(x,\cdot)\geq a_{x})>0$. Note that when $F(x,\omega)\geq a_{x}$, we have $\kappa(x,\omega,t)=1$ and $F^{\kappa(x,\omega,t)-1}(x,\omega,t)=F(x,\omega)$. Then by the fact that both $\mathbb{Q}_{x}(F(x,\cdot)\in I_{x})>0$ and $\mathbb{Q}_{x}(F(x,\cdot)\in J_{x})>0$, as well as $I_{x}$ and $J_{x}$ are disjoint, $F^{\kappa(x,\cdot,t)-1}(x,\cdot,t)$ can not be $\mathbb{Q}_{x,t}$-a.s. a constant.
This contradiction implies ¯$\overline{\mathbb{Q}}(G<\infty)<1$, or in other words, for $\theta_{E}$-positive $(x,t)\in E$, for $\mathbb{Q}_{x,t}$-positive $\omega\in\Omega$,
$\limsup_{n\to\infty}\overline{S}_{n}F(x,\omega,t)=+\infty.$
On the other hand, note that by construction, for $(x,t)\in E=X\times[-c,c]$, when $T_{f}^{E}(x,t)$ returns to $E$, we have
$t+f(x)+\cdots+f\circ T^{\rho_{E}(x,t)-1}(x)\in[-c,c].$
Thus for all $n\geq 1$,
$|\overline{S}_{n}f(x,\omega,t)|\leq 2c.$
Since $\theta_{E}=\mu\times\lambda_{c}$, we obtain that for $\mu$-positive $x\in X$, for $\mathbb{Q}_{x}$-positive $\omega$, there exists a subsequence $n_{k}=n_{k}(x,\omega)\to\infty$ such that $S_{n_{k}}f(x)\in[-2c,2c]$ for all $k\geq 1$ and $\limsup_{k\to\infty}S_{n_{k}}F(x,\omega)=+\infty$, therefore
$\limsup_{n\to\infty}S_{n}f(x)+S_{n}F(x,\omega)=+\infty.$

### v2

@@ 1. Introduction
is independent under $\mathbb{Q}_{x}$, therefore in this setting we have the following corollary from Theorem 1.3 (for simplicity we assume $h$ to be in $L^{1}$ and $\mathbb{Q}$ is ergodic).
Corollary 1.7.
Assume that
  • for $\mu$-a.e. $x\in X$, $h(x,\cdot)$ is not $\Theta_{x}$-a.s. a constant;
  • $h\in L^{1}(\Theta)$, $\int_{X\times Y}h\,\mathrm{d}\Theta=0$ and $\mathbb{Q}$ is $\hat{T}$-ergodic.
Then for $\mu$-a.e. $x\in X$ and for $\prod_{k=1}^{\infty}\Theta_{T^{k-1}(x)}$-a.e. $(y_{k})_{k\in\mathbb{N}}\in\bigotimes_{\mathbb{N}}Y$,
$\liminf_{n\to\infty}\sum_{k=1}^{n}h(T^{k-1}(x),y_{k})=-\infty\text{ and }\limsup_{n\to\infty}\sum_{k=1}^{n}h(T^{k-1}(x),y_{k})=+\infty.$
2. Proof of Theorem 1.3
We shall use the filling scheme, see [5] for example.
For $a\in\mathbb{R}$ denote by $a^{+}=\max\{a,0\}$ and $a^{-}=\max\{-a,0\}$.
For $n\geq 1$ define
$G_{n}:=\max_{1\leq k\leq n}\hat{S}_{k}F.$
We have
$G_{n+1}^{+}-G_{n+1}^{-}=G_{n+1}=F+G_{n}^{+}\circ\hat{T}.$
Let $G=\lim_{n\to\infty}G_{n}$. Then
(2.1) $G^{+}-G^{-}=F+G^{+}\circ\hat{T}.$
From the fact
$\mathbb{Q}\Big{(}\lim_{n\to\infty}\frac{\hat{S}_{n}F}{n}=0\Big{)}=1$
we deduce that, $\mathbb{Q}$-a.s., $\hat{S}_{n}F$ changes signs in the wide sense infinite many times, that is, $\hat{S}_{n}F>0$ or $\hat{S}_{n}F<0$ cannot hold for all $n$ large enough. This is due to a result of Kesten [6] that the sums of stationary sequences cannot grow slower than linearly, see [5, Section 5.(c)] for example. This implies that, $\mathbb{Q}$-a.s., $G\geq 0$ hence $G^{-}=0$, thus we may write from (2.1) that, $\mathbb{Q}$-a.s.,
(2.2) $G^{+}=F+G^{+}\circ\hat{T}.$
Assume that $\mathbb{Q}(G^{+}<\infty)=1$. This implies that for $\mu$-a.e. $x\in X$,
$\mathbb{Q}_{x}(G^{+}(x,\cdot)<\infty)=1.$
Fix $t\in\mathbb{R}\setminus\{0\}$. For $\mu$-a.e. $x\in X$ denote by
$\Phi_{t}(x):=\mathbb{E}_{\mathbb{Q}_{x}}(e^{itG^{+}(x,\cdot)})\text{ and }\phi_{t}(x):=\mathbb{E}_{\mathbb{Q}_{x}}(e^{itF(x,\cdot)}).$
Then by (2.2) and (A)(ii) we have, for $\mu$-a.e. $x\in X$,
$\Phi_{t}(x)=\phi_{t}(x)\times\Phi_{t}\circ T(x).$
Write $u(x)=-\log|\Phi_{t}(x)|$ and $v(x)=-\log|\phi_{t}(x)|$ with the convention that $-\log 0=\infty$. Then we have, $\mu$-a.s., $u,v\geq 0$ and
$u=v+u\circ T.$
Suppose that $\mu(u<\infty)>0$. Take $N>0$ large enough such that
$X_{N}=X\cap\{u\leq N\}$
has positive $\mu$-mass. Let $(X_{N},\mathcal{B}_{N},T_{N},\mu_{N})$ denote the induced dynamical system of $(X,\mathcal{B},T,\mu)$ by $X_{N}$, where $\mathcal{B}_{N}$ is $\mathcal{B}$ restricted to $X_{N}$, $\mu_{N}=\frac{1}{\mu(X_{N})}\mu|_{X_{N}}$ is the normalised $\mu$ restricted to $X_{N}$, and, denoting $\rho_{N}(x):=\inf\{n\geq 1:T^{n}(x)\in X_{N}\}$ the first visiting time to $X_{N}$, $T_{N}(x)=T^{\rho_{X_{N}}(x)}(x)$ is the induced transform from $X_{N}$ to $X_{N}$.
On $(X_{N},\mathcal{B}_{N},T_{N},\mu_{N})$ we obtain the co-boundary equation
$u-u\circ T_{N}=v_{N},$
@@
$v_{N}(x)=v(x)+v\circ T(x)+\cdots+v\circ T^{\rho_{X_{N}}(x)-1}(x).$
@@
On $X_{N}$ we have $0\leq u\leq N$, $0\leq u\circ T_{N}\leq N$ and $0\leq v_{N}\leq u\leq N$ are all bounded functions hence integrable. Thus by the $T_{N}$-invariance of $\mu_{N}$ and Birkhorff ergodic theorem we deduce that
$\mathbb{E}_{\mu_{N}}(v_{N}\,|\,\mathcal{I}_{N})(x)=0$
@@
for $\mu_{N}$-a.e. $x\in X_{N}$, where $\mathcal{I}_{N}$ is the $T_{N}$-invariant $\sigma$-algebra of $X_{N}$. Since $v$ is non-negative, $v_{N}$ is also nonnegative, we deduce that $v_{N}(x)=0$ and therefore $v(x)=0$ for $\mu_{N}$-a.e. $x\in X_{N}$. Since $N$ is arbitrary, we deduce that $v(x)=0$ or equivalently $|\phi_{t}(x)|=1$ for $\mu$-a.e. $x\in X\cap\{u<\infty\}$. Note that $|\phi_{t}(x)|=1$ implies $F(x,\cdot)$ is $\mathbb{Q}_{x}$-a.s. a constant, which is contradict to (A)(i). Therefore $\mu(u<\infty)=0$, or in other words,
(2.3) $\mu(x\in X:\Phi_{t}(x)=0)=1.$
We may take a countable sequence $t_{n}$ tending to 00. Then (2.3) implies that
$\mu(x\in X:\Phi_{t_{n}}(x)=0\text{ for all }n\geq 1)=1.$
But this is not possible since for $\mu$-a.e. $x\in X$, $\mathbb{Q}_{x}(G^{+}(x,\cdot)<\infty)=1$, that is, $G^{+}(x,\cdot)$ is a proper random variable under $\mathbb{Q}_{x}$, thus $\mathbb{E}_{\mathbb{Q}_{x}}(e^{itG^{+}(x,\cdot)})$ is continuous in $t$, hence non-vanishing around 00.
We finally deduce that $\mathbb{Q}(G^{+}<\infty)<1$, or in other words,
$\mathbb{Q}\Big{(}\limsup_{n\to\infty}\hat{S}_{n}F=+\infty\Big{)}>0.$
Since same arguments apply to $-F$, we also get
$\mathbb{Q}\Big{(}\liminf_{n\to\infty}\hat{S}_{n}F=-\infty\Big{)}>0.$